OpenStudy (anonymous):

(3x + 8) (x - 7) How do I solve this?

OpenStudy (anonymous):

3x(x-7) +8(x-7)

OpenStudy (anonymous):

\[(3x + 8) (x - 7)=3x^2-21x+8x-56 = 3x^2 -13x -56\]

OpenStudy (anonymous):

\[x ( 3x - 13) -56\]

OpenStudy (anonymous):

\[x=7 , x= -\frac{8}{3}\]

OpenStudy (imtant):

You need to expand it..Then factorize to get x values..There will be two values of x..

OpenStudy (imtant):

3x^2-13x-56=0 (x-7)(3x+8)..Then you will get the x values..Ok?

Directrix (directrix):

@JamesR4494 What are the instructions for this problem? Multiply and then simplify? Or, solve for x if (3x + 8) (x - 7) = 0 ? >>>>(3x + 8) (x - 7) How do I solve this?